### v1

# 1]Univ. Artois, UR 4025, Laboratoire Systèmes Electrotechniques et Environnement (LSEE), F-62400 Béthune, France 2]Univ. Artois, UR 3926, Laboratoire de Génie Informatique et d’Automatique de l’Artois (LGI2A), F-62400 Béthune, France Time series classification with random convolution kernels based transforms: pooling operators and input representations matter

Abstract
This article presents a new approach based on MiniRocket, called SelF-Rocket, for fast time series classification (TSC). Unlike existing approaches based on random convolution kernels, it dynamically selects the best couple of input representations and pooling operator during the training process. SelF-Rocket achieves state-of-the-art accuracy on the University of California Riverside (UCR) TSC benchmark datasets.
keywords:
Time series classification, ROCKET, MiniRocket, Pooling operators, Input representation, Feature selection.1 Introduction
Random convolution kernels based transforms, ROCKET [1] and its successors, MiniRocket [2], MultiRocket [3] and others, have revolutionized the field of time series classification, offering an unrivaled compromise between speed and accuracy [4].
The main idea behind ROCKET methods is to use a very large number of random convolution kernels to transform the time series and apply one or more pooling operators to generate a set of features. A simple linear classifier can then be used.
To the best of our knowledge, all the proposed transformations rely on one or more fixed pooling operators such as, for example, PPV (Proportion of positive values), for feature extraction from time series.
In this article, a new approach based on MiniRocket, called SelF-Rocket, is developed. Unlike existing ones based on random convolution kernels, it dynamically selects the best data input representations and pooling operator combination during the training process.
This article is organized as follows. The fundamentals of time series classification and the main random convolution kernels based methods are recalled in Section 2. A statistical analysis, conducted on the University of California Riverside (UCR) archive [5], is then presented in Section 3 to motivate this work. It shows that while PPV is a pooling operator of choice, in most cases it is not the best one. The method SelF-Rocket is then introduced in Section 4 and its performances on the UCR archive are analyzed in Section 5. Finally, Section 6 offers a conclusion with some perspectives.
2 Related works
2.1 Time series classification
Time series classification (TSC) is a type of supervised learning problem where the objective is to assign a class label to a given time series instance. A Time series data consist of sequences of observations collected at constant intervals over time. Formally, a time series can be represented as:
where denotes the value of the time series at time step , and represents the length of the time series. This notation will be used to refer to a time series throughout this article.
Many methods of various kind have been developed to address TSC problems. Traditional methods include distance-based techniques like Dynamic Time Warping (DTW) [6], which measure the similarity between time series. More recently, Deep Learning approaches such as Inceptiontime [7] have been proposed. Hybrid approaches like HIVE-COTE v2.0, combining multiple classifiers of different kind, are generally the most accurate, at a high computing cost [8]. For a recent general survey of TSC methods, the interested reader may refer to [4].
In the next section, we will focus on a specific type of TSC methods using random convolution kernels to extract features from time series data.
2.2 Random convolution kernels based methods
As mentioned above, these methods make use of random convolutions to discriminate time series. Convolution kernels can be seen as mini-random sequences whose convolution with time series generate activation maps used to extract several synthetic features that are subsequently fed into a simple linear classifier. The training phase thus only involves learning the appropriate weights assigned to each feature, which allows these methods to be, at the time this article is written, among the fastest while ensuring very good performance [4].
2.2.1 Main methods
ROCKET (Random Convolutional Kernel Transform) [1] is the first algorithm of this kind to have been introduced. It randomly generates a large number of convolution kernels (typically 10,000) which it uses to create activation maps. These maps are then summarized by two pooling operators: PPV (Proportion of Positive Values), which computes the percentage of positive values, and GMP (Global Maximum Pooling), which extracts the maximum value from the activation map. Formally:
where is the convolution output of .
For each time series, features are extracted, where is the number of kernels.
Although random, the kernels are parameterized as follows:
Length is randomly chosen from with equal probability.
Weights are randomly selected and then normalized, .
A bias is added to the activation map.
Dilations are computed with , where and , where is the length of the input time series, and the length of the kernel.
MiniRocket [2] is a variant of ROCKET with several key modifications that greatly speed-up the training without sacrificing performances. The method becomes much more deterministic by making the following changes:
Kernel length is fixed at 9 instead of .
It uses a single set of 84 kernels containing only as values. This major change optimizes convolution operations and significantly reduces computation time.
The bias is now derived from the convolution result with a kernel/dilation pair.
GMP is no longer calculated, only PPV is used.
MultiRocket [3] extends MiniRocket by introducing several improvements:
First-order difference () between two time units, corresponding to the rate of change, is used as an additional input representation. Formally:
(4) -
In addition to PPV, the following features are extracted: Mean of Positive Values (MPV), Mean of Indices of Positive Values (MIPV) and Longest Stretch of Positive Values (LSPV), formally defined by:
(5) where is the subset of positive value in ,
(6) where indicates the indices of positive values and
HYDRA (Hybrid Dictionary-ROCKET Architecture) [9] combines aspects of dictionary-based methods and random convolutions. The general idea is to group all kernels into groups of kernels. For each time series in the dataset, convolution is performed with all kernels from a given group. At each point in the series, the kernel leading to the highest convolution value is selected. Thus, for each group, a histogram of maximum kernel responses is created, serving as features for the classifier. First-order difference is used here as well. The kernels are parameterized as follows:
Length is fixed at 9.
Weights are randomly selected.
No bias or pooling operator is used after convolution.
The features generated by HYDRA and MultiRocket can be concatenated leading to a more accurate classification [9].
2.2.2 Kernel pruning
The main drawback of methods based on random convolutions is the large number of kernels, which may prevent them from being integrated into memory-limited devices. Moreover, not all kernels contribute positively to model performance, resulting in unnecessary computational complexity [10].
Various approaches have therefore been proposed for pruning the kernels, keeping only those most important to the problem.
S-ROCKET [10] selects the most important kernels and prunes redundant ones as follows:
1.
Pre-training the original model at full capacity,
2.
Initializing a population of binary candidate state vectors where each vector element represents the active/inactive status of a kernel,
3.
Using a population-based optimization algorithm to evolve this population, aiming to minimize the number of active kernels while maximizing classification accuracy,
4.
Combining the total number of active kernels with the classification accuracy to determine the best state vector, which is then used to select the optimal kernels.
Authors show that this approach maintains the classification performance of the original model while significantly reducing computational demands.
POCKET [11] improves S-ROCKET by dividing the pruning problem in 2 stages:
1.
Applying dynamically varying penalties for group-level regularization to prune redundant kernels.
2.
Uses element-level regularization to refit a linear classifier with the remaining features.
Experimental results demonstrate that POCKET performs 11 times faster than S-ROCKET while being as accurate.
Detach-ROCKET [12] prunes kernels using Sequential Feature Detachment (SFD), a method to identify and remove non-essential features. SFD involves iterative detachment steps, where a fixed percentage of the current active features is removed at each step. Depending on the number of classes, one or several RIDGE classifiers are trained on the set of active features, and RIDGE regressions are made to learn coefficients associated with each feature allowing one to eliminate the least relevant features.
2.2.3 ROCKET ensembles
Another way of improving a ROCKET classifier is to consider an ensemble of ROCKET methods.
Arsenal [8] is a set of small convolution-based ROCKET classifiers that produces a better probability estimate than ROCKET. It was introduced as a replacement of ROCKET in HIVE-COTE v2.0. As an ensemble method, Arsenal uses majority vote to classify new cases.
FT-FVC [13] adds feature diversity to ROCKET using three transformations to raw data: Hilbert and first-order and second-order difference. These transformations generate feature vectors that are concatenated with the raw feature vector to form a three-view representation. These representations are merged through hard voting.
3 On the importance of choosing the right input representations and pooling operator
To understand how input representations and pooling operators impact the performances of random convolution kernels based transforms, we compared, on the 112 datasets of the UCR archive [5], original MiniRocket with modified versions using GMP, MPV, MIPV and LSPV instead of PPV as pooling operator and the first order difference as additional input representation, leading to possible transforms (detailed in Section 4.1). In this article, transforms using multiple pooling operators were not considered since preliminary results showed they under-performed those using only one pooling operator in almost all the cases. Table 1 shows the mean classification accuracy across resamples of the possible transforms for some datasets included in UCR archive. For each dataset, the number in bold indicates the greatest overall accuracy achieved. Note that the PPV column is also the performance of the original MiniRocket classifier. Figure 1 shows the number of datasets in the UCR archive for which the modified MiniRocket version performs the best. If PPV_MIX is the best transform on average, it is outperformed by others in most cases ().
Figure 2 shows the critical difference diagram [14] of the 15 possible transforms for the different datasets, the distribution of classification accuracy is provided as box plots in Figure 13.
| DATASET | PPV | GMP | MPV | MIPV | LSPV | PPV | GMP | MPV | MIPV | LSPV | PPV | GMP | MPV | MIPV | LSPV |
| DIFF | DIFF | DIFF | DIFF | DIFF | MIX | MIX | MIX | MIX | MIX | ||||||
| ACSF1 | 82.77 | 55 | 82.97 | 69.23 | 85.27 | 79.9 | 60.6 | 81.8 | 68.6 | 84.4 | 82.2 | 61.4 | 83.03 | 70.47 | 85.57 |
| Adiac | 80.26 | 70.92 | 80.14 | 78.78 | 79.94 | 82.19 | 64.88 | 79.68 | 75.44 | 79.52 | 82.86 | 71.62 | 82.05 | 79.09 | 81.52 |
| ArrowHead | 88.15 | 86.34 | 87.6 | 85.85 | 88.91 | 87.9 | 83.41 | 85.77 | 81.22 | 88.57 | 88.7 | 86.67 | 88.06 | 85.05 | 89.49 |
| Beef | 77 | 61.44 | 75.22 | 73.56 | 68.67 | 77.78 | 66.22 | 71.56 | 77.78 | 67.56 | 79.67 | 68.89 | 74.56 | 77.44 | 70.56 |
| BeetleFly | 91 | 83 | 85.67 | 88 | 86.83 | 90.83 | 91.5 | 90.67 | 88.83 | 87.17 | 91.83 | 87.67 | 88.5 | 91 | 87.83 |
| BirdChicken | 92 | 90.33 | 89.83 | 83.67 | 91.33 | 90.67 | 93.33 | 91.83 | 88 | 89.83 | 90.17 | 91.33 | 88.5 | 85.33 | 90.17 |
| BME | 99.18 | 98.16 | 99.73 | 99.31 | 92.73 | 98.13 | 96.4 | 99.78 | 99.29 | 92.09 | 99.2 | 98.4 | 99.87 | 99.47 | 93.96 |
| Car | 91.78 | 87.28 | 91.44 | 91.61 | 88.83 | 91.06 | 76 | 89.17 | 88.89 | 83.94 | 92.61 | 86.33 | 92.33 | 92 | 89.33 |
| CBF | 99.62 | 98.7 | 99.4 | 96.09 | 99.39 | 83.23 | 76.34 | 80.93 | 79.71 | 81.63 | 99.36 | 97.69 | 99.24 | 94.38 | 98.88 |
| Chinatown | 96.87 | 95.17 | 96.66 | 97.28 | 96.14 | 95.83 | 90.39 | 94.88 | 96.36 | 94.7 | 96.56 | 94.19 | 96.03 | 97.14 | 95.96 |
| ChlorineConcent. | 75.42 | 73.45 | 71.75 | 75.86 | 73.29 | 78.69 | 82.53 | 78.07 | 78.55 | 75.64 | 78.15 | 81.86 | 77.59 | 78.65 | 75.81 |
| CinCECGTorso | 87.58 | 80.48 | 87.02 | 93.7 | 87.27 | 95.79 | 75.79 | 92.83 | 96.59 | 86.58 | 92.56 | 80.79 | 90.74 | 96.09 | 89.56 |
| Coffee | 99.88 | 99.88 | 99.88 | 100 | 100 | 98.69 | 97.26 | 99.88 | 99.64 | 98.57 | 99.76 | 99.29 | 100 | 100 | 100 |
| Computers | 80.16 | 73.12 | 81.44 | 72.89 | 79.36 | 84.84 | 67.35 | 82.23 | 74.96 | 79.95 | 84.87 | 73.87 | 84.56 | 75.75 | 82.6 |
| CricketX | 82.62 | 73.26 | 81.25 | 79.51 | 80.77 | 70.76 | 50.56 | 68.62 | 65.49 | 69.4 | 81.74 | 74.19 | 81.86 | 78.2 | 79.74 |
| CricketY | 84.24 | 74.61 | 82.51 | 79.46 | 81.45 | 68.76 | 45.69 | 67.48 | 65.8 | 68.08 | 83.51 | 75.06 | 82.92 | 80.08 | 80.78 |
| CricketZ | 84.25 | 75.86 | 83.42 | 81.2 | 82.44 | 72.18 | 49.69 | 70.23 | 65.42 | 71.68 | 83.14 | 76.36 | 83.69 | 79.66 | 82.17 |
| Crop | 76.37 | 69.89 | 76.17 | 75.94 | 76.25 | 70.65 | 63.26 | 71.04 | 70.66 | 70.25 | 76.71 | 71.86 | 76.99 | 76.12 | 76.37 |
| DiatomSizeReduction | 94.29 | 96.21 | 94.61 | 95.68 | 95.46 | 93.69 | 92 | 94.07 | 95.46 | 92.02 | 94.23 | 95.41 | 94.62 | 95.92 | 95.17 |
| DistalPha.Out.AgeGr. | 79.42 | 78.13 | 79.66 | 80.12 | 79.62 | 80.17 | 77.31 | 81.03 | 80.41 | 79.57 | 79.74 | 77.53 | 79.62 | 80.05 | 80.19 |
| DistalPha.Out.Corr. | 82.44 | 79.83 | 82.57 | 82.95 | 82.73 | 83.97 | 80.92 | 83.32 | 83.85 | 82.8 | 83.93 | 81.44 | 83.67 | 83.8 | 83.47 |
| DistalPhalanxTW | 69.5 | 67.84 | 69.81 | 69.47 | 69.06 | 69.9 | 67.39 | 69.95 | 68.25 | 68.9 | 69.81 | 67.63 | 69.66 | 68.85 | 68.94 |
| Earthquakes | 73.79 | 73.45 | 73.93 | 74.96 | 73.96 | 74.6 | 71.7 | 74.2 | 74.82 | 74.82 | 74.12 | 73.74 | 74.15 | 75.2 | 74.44 |
| ECG200 | 90.2 | 88.57 | 89.67 | 88.43 | 89.17 | 84.63 | 83.17 | 85.07 | 83.9 | 82.43 | 89.67 | 86.07 | 89.17 | 87.6 | 87.77 |
| ECG5000 | 94.65 | 94.64 | 94.74 | 94.56 | 94.53 | 94.21 | 94.03 | 94.41 | 94.26 | 94.18 | 94.65 | 94.64 | 94.74 | 94.57 | 94.6 |
| ECGFiveDays | 99.07 | 99.69 | 99.65 | 97.5 | 99.55 | 98.83 | 95.74 | 99.04 | 96.88 | 98.35 | 99.21 | 99.77 | 99.73 | 97.23 | 99.66 |
| ElectricDevices | 87.44 | 74.46 | 87.71 | 85.12 | 87.07 | 85.83 | 69.01 | 85.47 | 79.61 | 84.79 | 89.16 | 76.26 | 88.99 | 86 | 88.13 |
| EOGHorizontalSignal | 83.46 | 79.91 | 84.66 | 80.9 | 83.49 | 80.02 | 65.91 | 80.71 | 73.49 | 76.73 | 85.17 | 78.8 | 86.56 | 81.57 | 84.79 |
| EOGVerticalSignal | 79.91 | 76.02 | 80.9 | 76.77 | 79.24 | 73.78 | 58.79 | 75.78 | 68.69 | 73.27 | 79.4 | 74.19 | 82.14 | 76.57 | 79.25 |
| EthanolLevel | 66.27 | 58.43 | 62.74 | 71.87 | 70.86 | 60.56 | 37.51 | 51.85 | 65.53 | 51.55 | 66.75 | 56.09 | 62.42 | 71.49 | 69.89 |
| FaceAll | 98.35 | 96.96 | 98.65 | 98.31 | 97.96 | 96.88 | 93.2 | 97.07 | 97.11 | 96 | 98.36 | 96.91 | 98.64 | 98.26 | 97.81 |
| FaceFour | 94.24 | 81.25 | 92.31 | 90.08 | 92.99 | 78.03 | 59.39 | 75.3 | 81.14 | 78.56 | 92.8 | 77.58 | 89.62 | 88.64 | 90.38 |
| FacesUCR | 96.92 | 93.5 | 96.71 | 97.11 | 95.87 | 92.4 | 86.88 | 93.13 | 93.54 | 90.93 | 96.6 | 93.57 | 96.63 | 96.77 | 95.2 |
| FiftyWords | 82.89 | 72.59 | 81.53 | 82.32 | 79.71 | 80.13 | 53.93 | 74.44 | 79.42 | 74.63 | 83.69 | 72.64 | 81.25 | 82.31 | 79.68 |
4 SelF-Rocket
Following the experimental results obtained in the previous section, we investigate a new variant of ROCKET, called SelF-Rocket (Selected Features Rocket), that aims at dynamically selecting the best couple of input representations and pooling operator during the training process. The latter involves three main stages:
SelF-Rocket works in a similar way to MiniRocket or MultiRocket. The only major difference between them is the addition of a feature selection phase. As seen previously in Section 3, the optimal set of features varies from one dataset to another, so using a single feature type (PPV) like MiniRocket or a fixed set of feature types ({PPV, LSPV, MIPV, MPV}) like MultiRocket can produce good results on average, but not optimal for each case. To address this problem, we implement a wrapper-based feature selection module with the objective to select the most appropriate set for each classification problem.
4.1 Feature Generation
Let IR be a set of Input Representations, PO a set of Pooling Operators, MK the set of all kernels (containing every kernel/dilation combinations) generated by MiniRocket and the bias associated to a kernel (cf Section 2.2.1).
In the implementation evaluated in this article we consider
and
where is the identity function ( for all time series ) and is the first order difference (cf Equation 4).
Let be a feature generation function defined as:
where is a time series, a subset of the power set of IR (minus the empty set), i.e. , and a pooling operator, i.e. .
Thus, following this formalization, the original MiniRocket extracts features using the parametrization , for any time series .
The Cartesian product induces a set of such parameterized functions; In this implementation, and , leading to possible parametrizations.
Let FV be the generated Feature Vector. In the following, feature generation outputs are denoted as in Section 3 for readability reasons:
p for ,
p_DIFF for ,
and p_MIX for ,
for all .
4.2 Feature Selection
The Feature Selection Module relies on the Stratified -Fold method to create novel train and validation sets derived from the original train set. These sets are then used to train mini-classifiers for each IR-PO combination. Subsequently, the optimal combination on average is selected as the final set of features for the linear classifier. The Feature Selection Module is summarized in Algorithm 1.
Training set class y_train
The Number of folds k
The Number of features per mini-classifier f
4.3 Classification
SelF-Rocket employs the same algorithm for the mini-classifiers embedded within the Feature Selection Module and the end-stage classifier, namely the RIDGE classifier, as proposed by Dempster et al. [1]. This classifier is preferable to stochastic descent methods, such as logistic regression, when the number of features exceeds the number of training examples and when working with small datasets, e.g. when there are fewer than 10,000 examples.
5 Experiments
In this section, the performances of SelF-Rocket are evaluated.
We show that SelF-Rocket is as accurate as Hydra-MultiRocket and HIVE-COTE v2.0, despite having only one type of feature, and also has a relatively shorter classifier training time. We also investigate the
influence of the main parameters on the model performance.
5.1 Experimental settings
SelF-Rocket has been evaluated on a subset of the UCR Time Series Archive [5], which is widely used in this field. To guarantee the reproducibility of our experiments and a fair comparison with other TSC algorithms, we employed the identical 112 datasets as those used by [4] and the precise same 30 train/test resamples for each of those datasets. The 112 datasets are those from the 128 original ones that did not contain missing values or variable time series length.
Note that we did not compute the results for others methods shown in Figures 4, 5 and 6, but used the results111https://github.com/time-series-machine-learning/tsml-eval/tree/main/results/classification/Univariate available in [4].
The author’s implementation of MiniRocket222https://github.com/angus924/minirocket/tree/main was used as baseline model. Our algorithm is implemented in standard python 3.11.4.
In order to compare the classification performance with other TSC algorithms, we use the critical difference diagrams that display mean ranks of each method across all datasets with the horizontal cliques indicating that there is no statistically significant difference between those methods, the Multiple Comparison Matrix (MCM) [15] with heat-map color representing mean differences in score, and the pairwise scatter plots of test accuracy summarizing the number of win/draw/loss between 2 classifiers across all datasets.
All experiments were conducted on a Dell Precision 5480 on Debian 12 with an Intel Core i9-13900H vPro 2.6 GHz and 32GB of RAM
5.2 Comparison with other TSC methods
We compare the performances of SelF-Rocket with 11 TSC methods: 1NN-DTW, Elastic Ensemble (EE) [16], Shapelet transform classifier (STC) [17], HIVE-COTE v2.0 [8], H-Inceptiontime [18], MultiRocket [3], Hydra, Hydra+MultiRocket [9], MiniRocket [2],
ROCKET [1] and QUANT [19].
Figure 4 shows the average mean accuracy rank of these classifiers. In terms of performance, SelF-Rocket ranks second only to HIVE-COTE v2.0, and outperforms others.
Figure 5 shows the Multiple Comparison Matrix [20], i.e. the full pairwise comparison between the best performing methods tested. The mean difference in accuracy between SelF-Rocket and its baseline MiniRocket is approximately 1.23 %
Figure 6 displays the mean accuracy, the pairwise win/draw/loss, and value for statistical tests between the tested methods over 30 resamples for the 112 selected UCR datasets. Each point represents a dataset, the more a point is distant to the diagonal line, the more one of the two methods performs better than the other. In some cases, SelF-Rocket outperforms Hydra-MultiRocket, but in general, both methods offer comparable performance. HIVE-COTE v2.0 also has the advantage over SelF-Rocket in certain datasets.
There is no significant statistical difference by Wilcoxon signed rank test between SelF-Rocket, Hydra-MultiRocket and HIVE-COTE v2.0
SelF-Rocket (9,996 or 19,992 features) is faster in the training and prediction phase than MultiRocket and Hydra-MultiRocket due to a lower number of features (resp 50,000 and 50,000 + number of features of Hydra).
5.3 Sensitivity Analysis and Ablative Study
As previously shown, SelF-Rocket outperforms other existing methods based on random convolution kernels. In this section, the impact of the three key parameters on global accuracy performance is studied:
Following [21, 19], we choose 50 ‘development’ datasets from the 112 UCR datasets to conduct this analysis. We first remove the 12 datasets with the most training examples, then we select the sets that contain at least 5 examples per class and 100 training examples. From the remaining list, 50 were randomly selected. The complete list is available in Table 4.
The others parameters of SelF-Rocket, i.e. number of kernels, padding, dilation, bias, length of kernel, values inside the kernel, remain the same as the default parameters of MiniRocket.
5.3.1 Number of folds
The SelF-Rocket Feature Selection module employs a stratified -fold decomposition of the training set. We have conducted tests using the minimal number of folds possible, namely , up to a fixed value of . Figure 7 illustrates the impact of the number of folds on the classifier mean accuracy. No notable differences were observed when different values of were used.
5.3.2 Number of features
In its default configuration, MiniRocket only generates 9,996 kernels using PPV as pooling operator, whereas SelF-Rocket employs 9,996 kernels with the base input representation and an additional 9,996 with the first-order difference input representation.
For each couple (input representation, pooling operator) belonging to , SelF-Rocket generates 9,996 features. The number of features tested ranges from 2,500 to 9,996 for each mini-classifier within the Feature Selection module, features are randomly selected from the total number of features, which is either 9,996 or 19,992 ().
Figure 7 shows that the number of features has a relatively minor impact on the average accuracy of the final classifier.
5.3.3 Input Representations and Pooling Operators
| SelF-Rocket | SelF-Rocket | SelF-Rocket | SelF-Rocket | |
| ONLY BASE | ONLY DIFF | ONLY MIX | ||
| Win | – | 2 | 4 | 6 |
| Draw | – | 18 | 15 | 23 |
| Lose | – | 30 | 31 | 21 |
| Mean accuracy | 85.69 | 85.01 | 82.13 | 85.35 |
We evaluated the classification performance of the features generated by 3 sets of input representations: the raw time series (}), the first-order difference applied to the time series (}), and a concatenation of features for both (}). Figure 8 illustrates the discrepancy in ranking between the utilization of all three sets of representations and the exclusive use of a single one.
Table 2 provides a summary of the pairwise win/draw/loss between SelF-Rocket () and SelF-Rocket with only one set of representations.
The first-order difference is generally the least impactful representation, while the MIX set of representations demonstrates the best overall performance among the three.
| SelF-Rocket | SelF-Rocket | SelF-Rocket | SelF-Rocket | SelF-Rocket | SelF-Rocket | |
| NO PPV | NO GMP | NO MIPV | NO MPV | NO LSPV | ||
| Win | – | 4 | 1 | 1 | 0 | 4 |
| Draw | – | 27 | 48 | 39 | 46 | 43 |
| Lose | – | 19 | 1 | 10 | 4 | 3 |
| Mean acc. | 85.69 | 85.44 | 85.63 | 85.36 | 85.60 | 85.67 |
The choice of a pooling operator or a set of pooling operators to generate features can affect the performance of the classifier in the classification process. Figure 9 and Table 3 show the impact on performance of not using one of the five pooling operators. It can be seen that not using GMP or LSPV has no significant impact on system performance. On the other hand, not using MIPV or PPV has a greater impact on overall performance.
5.4 Scalability and Compute Time Comparison
Compared to other ROCKET methods, SelF-Rocket incorporates a Feature Selection module between the global set of features and the final classifier. The time required for the generation of SelF-Rocket features is less than Hydra-MultiRocket, while the training of the classifier is similar in duration to that of MiniRocket.
Figure 10 shows the mean compute time for all resamples and all datasets of the Feature Selection module according to the number of folds and features. The classifier complexity depends on , the number of features and , the number of training examples. is a multiplier for the number of mini-classifiers, and impacts the complexity of a single classifier so the two parameters impact generally the compute time.
For each combination of IR and PO, we instantiate mini-classifiers to identify the optimal set of features. A RIDGE classifier with a complexity of is employed. In Figure 11(a) given that is a fixed value, namely 9,996, we can see that the mean time only depends on the number of training examples.
Figure 12(a) shows the average total computation time over the 50 development datasets, for all resamples. SelF-Rocket and MiniRocket training (Figure 12(a)) and prediction times (Figure 12(b)) are comparable. However, among the evaluated methods, MiniRocket remains the fastest overall. As illustrated in Figure 10, with and SelF-Rocket’s average feature selection time is around 1.3 seconds, which is significantly faster than its feature creation computation time. SelF-Rocket features take longer to create than MultiRocket features. This is due to the greater number of kernels that are instantiated. MultiRocket instantiates only 6,216 kernels for each of its input representations (DIFF and BASE). From these, features () are generated. In comparison, SelF-Rocket generates 9,996 kernels for each of these two representation types for a total of features ().
SelF-Rocket produces times more features than MiniRocket. One way to improve its overall compute time could be to reduce the number of kernels generated by SelF-Rocket or to reduce the number of input representations and pooling operators.
HYDRA produces kernels : (number of kernels per group), (number of groups) and (maximum possible dilation value), with 2 features extracted per kernel (maximum and minimum responses). For each dataset, 512 ( and ) kernels per dilation are created with . In most of our cases will be between 7 and 10, so the total number of features generated will generally be between (7,128) and (10,240). This explains why HYDRA is the fastest in terms of classifier training and testing time. As illustrated in Figures 12(c) and 12(d), SelF-Rocket is faster than Hydra-MultiRocket in terms of both feature creation time and classifier testing time, with less discrepancy.
6 Conclusion and prospects
The two main contributions of this article are as follows:
First, we have shown that the choice of Input Representations and Pooling Operators significantly impacts the performance of TSC methods based on random convolution kernels. We have also shown that using multiple Pooling Operators, such as MultiRocket, leads to good performances on average, but is not optimal in specific cases: it is always preferable to select a single appropriate Pooling Operator.
Then, on this basis, we have proposed a new algorithm based on MiniRocket, SelF-Rocket, incorporating a feature selection stage. Experiments show that this algorithm selects the best Input Representations / Pooling Operator combination in most cases, leading to state of the art performances. In the tested implementation of SelF-Rocket, only 9,996 or 19,992 features are retained to train the RIDGE classifier, far fewer than competing methods such as MultiRocket or Hydra-MultiRocket, leading to fast predictions.
In this article, a minimal version of SelF-Rocket was evaluated, using 2 input representations, 5 pooling operators and a simple wrapper-based feature selection method. This implementation could be improved by:
Considering more input representations (second order difference, Fourier and Hilbert transforms, etc.),
Combining SelF-Rocket and HYDRA features,
Using HYDRA as a pooling operator within SelF-Rocket,
Using a filter-based feature selection method (based on test for instance) instead of Stratified -Fold to speed-up the feature selection stage.
Acknowledgments
This research project, supported and financed by the French ANR (Agence Nationale pour la Recherche), is part of the Labcom (Laboratoire Commun) MYEL (MobilitY and Reliability of Electrical chain Lab) involving LSEE, LGI2A and CRITTM2A (ANR-22-LCV2-0001 MYEL).
Appendix A Additional Figures & Tables
| Developement datasets | ||
| ACSF1 | ChlorineConcentration | Computers |
| CricketX | CricketY | CricketZ |
| DistalPhalanxOutlineAgeGroup | DistalPhalanxOutlineCorrect | DistalPhalanxTW |
| Earthquakes | ECG200 | EOGHorizontalSignal |
| EOGVerticalSignal | EthanolLevel | FaceAll |
| Fish | FreezerRegularTrain | GunPointAgeSpan |
| GunPointMaleVersusFemale | GunPointOldVersusYoung | Ham |
| Haptics | InlineSkate | InsectWingbeatSound |
| LargeKitchenAppliances | MedicalImages | MiddlePhalanxOutlineAgeGroup |
| MiddlePhalanxOutlineCorrect | MixedShapesRegularTrain | MixedShapesSmallTrain |
| OSULeaf | Plane | PowerCons |
| ProximalPhalanxOutlineAgeGroup | ProximalPhalanxOutlineCorrect | ProximalPhalanxTW |
| ScreenType | SemgHandGenderCh2 | SemgHandMovementCh2 |
| SemgHandSubjectCh2 | ShapesAll | SmoothSubspace |
| SwedishLeaf | SyntheticControl | Trace |
| UWaveGestureLibraryX | UWaveGestureLibraryY | UWaveGestureLibraryZ |
| Worms | WormsTwoClass | |